OpenStudy (anonymous):

on june 1, 2009, the price of gas was $1.79 per gallon. The price is predicted to increase .3% per month. if that prediction is correct, about how many months, to the nearest whole number, will it take for the price of a gallon to reach $1.90?

OpenStudy (anonymous):

i neeed helpppp :[ trig regents tommarow!

OpenStudy (anonymous):

? anyone

OpenStudy (anonymous):

!!!!!

mathslover (mathslover):

can u rewrite the question please

OpenStudy (anonymous):

k

OpenStudy (anonymous):

on june 1, 2009, the price of gas was $1.79 per gallon. The price is predicted to increase .3% per month. if that prediction is correct, about how many months, to the nearest whole number, will it take for the price of a gallon to reach $1.90?

OpenStudy (anonymous):

A:1 B:2 C:19 D:20

OpenStudy (anonymous):

From your problem we have,\[p(m)=1.79\times(1+0.003)^{m}\] p being the price after m months.\[p(m)=1.90\rightarrow1.79\times(1+0.003)^{m}=1.90\rightarrow(1+0.003)^{m}=1.0614525\]Thus, \[m=19.909226\approx20\]So I would say that after 20 months you would have the oil price at 1.90.

mathslover (mathslover):

It will take 20 months. Since it is increasing 0.3% each month (that means that after each month the price increases, 0.3% of the previous month. i.e. a recursive calculation... You can either use a summation formula or (faster) to just keep multiplying your answer by 1.003 each time since it was a small number of price increases.

OpenStudy (anonymous):

okay.. thnks for the answer but that just confused me

mathslover (mathslover):

 Predicted Cost = (1.79) • (1.003^m) ... where m = months after June 1, 2009         1.90 = (1.79) • (1.003^m)    1.90 ⁄ 1.79 = 1.003^m   log[1.90 ⁄ 1.79] = m • log[1.003]           m = 20 months

OpenStudy (anonymous):

ooh so i wouldd jus plug in the answers n see which makes sense?

mathslover (mathslover):

k

OpenStudy (anonymous):

yup